OpenStudy (anonymous):

A quadrilateral PQRS is inscribed in a circle, as shown below. What is the measure of arc PQR? 260° 200° 100° 320°

OpenStudy (anonymous):

@david111 help again lol

OpenStudy (anonymous):

i have 4 more questions left after this

OpenStudy (anonymous):

oh boy.

OpenStudy (anonymous):

haha i know thats why i came here

OpenStudy (anonymous):

200 degrees

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem! whats the other ones?

OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y - 2)2 = 9

OpenStudy (anonymous):

(x + 1)2 + (y - 2)2 = 9

OpenStudy (anonymous):

do u want steps? lol

OpenStudy (anonymous):

no i just need to get done but thanks for offering im gonna open a new question so i can give you another medal

OpenStudy (anonymous):

ok ill give u metals too so u can get in my league